OpenStudy (anonymous):

write an equation for a linear function f that has the given values f(-2)=-2, f(4)=-8?

OpenStudy (hoblos):

the line pass by (-2,-2) & (4,-8) slope= -1 y+2=-(x+2) y=-x -4

OpenStudy (hoblos):

=> f(x) = -x-4